Paul lends R10 000 to his friend Steven for one year. They agree that Steven would pay the R10 000 back with 5% interest at the end of the year. If the inflation rate is 6% at the end of the year, which of the following would be true for the real value of the amount that Steven pays back ?


1
Expert's answer
2021-10-06T09:45:00-0400


Loan amount = R 10,000

Nominal rate = 5%

Inflation rate = 6%

Therefore

      "Real rate = [\frac{(1+Nominal \space rate)}{(1+Inflation\space rate)}]-1\\

 = [\frac{(1+0.05)}{(1+0.06)}]-1\\

 = \frac{1.05}{1.06}-1\\

 = 0.990566037735849-1\\

 = -0.00943396226415105"


As per formula

            Real value after 1 year = Loan amount"\times" (1+Real rate)

                                                   "= 10000\times(1-0.00943396226415105)\\

 = 10000\times 0.990566037735849\\

 = 9905.66"


Real value of the amount that Steven pays back = R 9905.66